OpenStudy (anonymous):

Consider the line that passes through the points (-5, 9) and (-2, 0). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 9 = -1 third(x + 5) y - 9 = -3(x + 5) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (anonymous):

some one can help me

OpenStudy (radar):

The equation of the line that would pass through those two given points would have a slope of -1/3. So only the first equation would apply.

OpenStudy (radar):

But let me double check. Yep.

OpenStudy (radar):

Do you want to see the step by step solution?

OpenStudy (anonymous):

ok it better and thank u

OpenStudy (radar):

slope M= (9-0)/(-5-(-2))=9/-3=1/-3

OpenStudy (radar):

Then the equation for that line would be: 1/-3=(y-9)/(x=7) cross multiply getting: -x-5=3y-27 rearranging you have 3y-27=-x-5 dividing both sides by 3 you get: y-9=-(1/3)(x+5) doesn't that look familiar? that is equation 1 and the first answer is correct.

OpenStudy (anonymous):

oh ok I get it now thank for you're time

OpenStudy (radar):

NP, good luck with your studies.

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

can you help me in other one